solve for x, and please show your work so I can figure out how to do it myself:
3(5)^(2x-3)=6

Answer by stanbon(75887)

3(5)^(2x-3)=6
---
Divide both sides by 3 to get:
5^(2x-3) = 2
---
Take the log of both sides:
---
(2x-3)log(5) = log(2)
---
2x-3 = log(2)/log(5)
-------
2x-3 = 0.4307
---
2x = 3.4307
x = 1.7154
=================
Cheers,
Stan H.
=================